OpenStudy (anonymous):

I need to convert 600 N/m^2 to a form that involves kg. 1 N = 1 kg*m/s^2 Not sure what I should do.

OpenStudy (whpalmer4):

Write out the quantity: 600 N/m^2 Now multiply by a fraction which will cancel out the unit you don't want. You want to cancel out N, so multiply by 1 (kg*m/s^2)/1 N \[600 \frac{N}{m^2} * \frac{1 kg*m/s^2}{1 N}=600 \frac{N}{m^2}*\frac{1 kg*m}{1 N*s^2}\] What are we left with after canceling common factors? Alternatively, because you have a 1:1 conversion factor, you could simply write "kg*m/s^2" in the first expression wherever you had N, and simplify that. \[600\frac{kg*m/s^2}{m^2} = 600 \frac{kg*m}{m^2s^2}\]